OpenStudy (anonymous):

Part 1: Find the area of a rectangular object which has a length of 2 radical 3 inches and a width of 3 radical 6 inches. (3 points) Part 2: Explain, in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (anonymous):

do you still need help or you good?

OpenStudy (anonymous):

The rules for multiplying radicals are as follows \[A \sqrt{B }* C \sqrt{D }= A*C \sqrt{B*D} \] Therefore \[area = 2\sqrt{3}*3\sqrt{6}\] \[A=6\sqrt{18} --------A=18\sqrt{3}\]

OpenStudy (anonymous):

sorry typo A= 18Sqrt(2)